# Hallucinate at the Last in Long Response Generation: A Case Study on Long Document Summarization

### v2

@@ Abstract
Large Language Models (LLMs) have significantly advanced text generation capabilities, including tasks like summarization, often producing coherent and fluent outputs. However, faithfulness to the source material remains a significant challenge due to the generation of hallucinations by LLMs. While extensive research focuses on detecting and reducing these inaccuracies, less attention has been paid to the positional distribution of hallucinations within generated text, particularly in long outputs. In this work, we investigate where hallucinations occur in LLM-based long response generation, using long document summarization as a key case study. Focusing on the challenging setting of long context-aware long response generation, we find a consistent and concerning phenomenon: hallucinations tend to concentrate disproportionately in the latter parts of the generated long response. To understand this bias, we explore potential contributing factors related to the dynamics of attention and decoding over long sequences. Furthermore, we investigate methods to mitigate this positional hallucination, aiming to improve faithfulness specifically in the concluding segments of long outputs.
@@ 1 Introduction
Figure 1: Comparison of the faithfulness scores of summaries generated around 50, 200, and 800 words in length for a 6K-token length Wikipedia context using GPT-4o mini.
@@ 2 Generating Summaries
This section describes the experimental setup used to generate the summaries analyzed in our study. We prompt LLMs to generate summaries from original documents, varying both the context length and the output length. For all experiments, we generate summaries using greedy decoding. The prompt is provided in Appendix A.
@@ 2.1 Input Context Lengths
To analyze the impact of input length, we evaluate model performance across a range of context lengths, varying from approximately 1K tokens up to 8K tokens. For the Llama model specifically, we further extend the evaluation to context lengths of 12K and 16K tokens.
@@ 2.2 Output Length Categories
This distinction allows us to compare hallucination patterns in standard summary lengths versus significantly longer generated responses. By including both standard and long output categories, we aim to systematically analyze if and how the positional bias of hallucinations manifests and potentially becomes more severe as the length of the generated summary increases. See Appendix A for more details.
@@ 2.3 Overall Faithfulness
As a preliminary step, we measure the overall faithfulness of the generated long summaries with FineSurE Song et al. (2024) to ensure consistent quality across different context lengths, prior to our main positional analysis. Table 1 shows results for Llama3.1-8B-Instruct Grattafiori et al. (2024) on the Wikipedia dataset (1K–8K context lengths), evaluated using FineSurE with GPT-4 Achiam et al. (2023). FineSurE assesses each sentence against the source, identifying error types when inconsistencies occur. The results align with those on CNNDM Hermann et al. (2015), where Llama3-70B-Instruct achieved 85.5% faithfulness. We next conduct a finer-grained positional analysis in Section 3 for various setups.
@@
Table 1: Faithfulness scores of long summaries on the Wikipedia dataset, evaluated using FineSurE.
3 Where Do Hallucinations Most Frequently Occur?
Figure 2: Comparison of faithfulness scores for standard summaries generated by different models across increasing output lengths on the Wikipedia dataset, with context lengths ranging from 1K to 8K.
@@ 3.1 Experimental Setup
We evaluate our approach on multiple state-of-the-art LLMs, including Llama3.1-8B-Instruct, Gemma3-12B-Instruct Team et al. (2025), Qwen2.5-7B-Instruct Yang et al. (2024a), and GPT-4o mini, to ensure model-agnostic findings. For Gemma, we retain the default sliding window attention context length of 1024 tokens, while for Qwen, we modify the default setting of over 131K tokens to 16K tokens.
@@
To examine whether the phenomenon is domain-specific, we analyze the positional distribution of faithfulness within generated summaries across various domains. In addition to the Wikipedia dataset used in the main experiment, we include the arXiv Cohan et al. (2018), PubMed Cohan et al. (2018), and GovReport Huang et al. (2021) datasets for further evaluation.
@@ 3.2 Evaluation Metric
To enable an in-depth analysis of positional bias within each generated summary, we use the MiniCheck-Flan-T5-Large Tang et al. (2024) model to evaluate faithfulness. MiniCheck is particularly suitable for this granular task, as it is an efficient and high-performing fact-checking model that leverages atomic facts Min et al. (2023); Jing et al. (2024); Yu et al. (2024).
Let $s_{i}$ be the $i$-th sentence in the generated summary $S$. Each sentence in the generated summary is decomposed into a set of atomic facts $A_{i}$ by utilizing LLM, followed by a filtering process to remove unnecessary atomic facts Yang et al. (2024b), which makes:
@@
Using a MiniCheck model, which is denoted as $\mathcal{M}$, each filtered atomic fact is then compared with the source document $D=\{d_{1},...,d_{M}\}$ in a pair-wise manner. Each atomic fact is assigned the highest score among its comparisons:
@@
Finally, we use the average of the scores for a set of filtered atomic facts in a sentence as its faithfulness score:
@@
Additional experimental details and an example of the evaluation procedure are provided in Appendix B.
Furthermore, we verify the quality of atomic facts generated by the LLM via human evaluation in Table 2 for atomic facts extracted from long summaries generated by Llama.
Context Length Validation(%)
1K 98.55
2K 99.08
3K 97.62
4K 95.92
5K 99.01
6K 96.18
7K 99.12
8K 99.64
Table 2: Human evaluation of the quality (completeness) of atomic facts. For each context length, we randomly sampled 10 summaries and conducted the evaluation manually. Each atomic fact was annotated as either true or false depending on whether it is present in the summary.
@@
In addition to the faithfulness score, we propose a simple yet effective metric for analyzing positional discrepancies in generated outputs. Specifically, following Singh and Goshtasbpour (2022), each summary is divided into five equal length bins or segments, and the average faithfulness score of the atomic facts within each bin is computed. We define sensitivity as the difference between the faithfulness score of the last bin and the average of the first four bins. A positive sensitivity value indicates a tendency to hallucinate at the last, whereas a negative value suggests otherwise. The larger the sensitivity, the more pronounced the hallucination at the end of the output.
@@
Llama3.1-8B-Instruct Overall Avg. 0.81 0.82 0.79 0.80 0.75 5.5
12K 0.80 0.79 0.66 0.63 0.69 9.5
@@
Llama3.1-70B-Instruct Overall Avg. 0.83 0.82 0.83 0.81 0.77 5.25
Gemma3-4B-Instruct Overall Avg. 0.85 0.82 0.83 0.85 0.79 4.75
Gemma3-12B-Instruct Overall Avg. 0.85 0.83 0.84 0.83 0.79 4.75
GPT-4o mini Overall Avg. 0.83 0.81 0.81 0.78 0.76 4.75
Qwen2.5-7B-Instruct Overall Avg. 0.86 0.83 0.83 0.82 0.84 -0.5
@@
Table 3: Faithfulness scores and sensitivity for long summaries generated by various models on the Wikipedia dataset. Overall Avg. denotes the average faithfulness score aggregated across context lengths from 1K to 8K.
3.3 Hallucinate at the Last
Model Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Llama3.1-8B-Instruct 1K 0.92 0.90 0.91 0.87 0.68 22.00
2K 0.82 0.74 0.83 0.72 0.71 6.75
3K 0.94 0.94 0.81 0.86 0.68 20.75
4K 0.92 0.92 0.88 0.87 0.82 7.75
5K 0.91 0.88 0.89 0.78 0.70 16.50
6K 0.86 0.84 0.79 0.89 0.73 11.50
7K 0.97 0.89 0.82 0.94 0.66 24.50
8K 0.93 0.90 0.86 0.79 0.65 22.00
@@ 3.3 Hallucinate at the Last
Table 4: Human evaluation of faithfulness scores and sensitivity for long summaries generated by Llama on the Wikipedia dataset. Each atomic fact was annotated as either true or false depending on whether it is present in the document. The results were manually evaluated and aggregated.
The latter part of the summary exhibits the lowest level of faithfulness across multiple models, consistently observed across various contexts and summary lengths. The decline becomes most pronounced when the context length reaches 8K and the summaries are long.
We present the positional faithfulness discrepancies of both standard and long summaries generated by different models on the Wikipedia dataset, with respect to varying input context lengths, in Figure 2 and Figure 3, respectively. As shown in Figure 2, standard summaries generated by the three models, excluding Qwen, consistently exhibit a decline in faithfulness towards the latter positions of the summary, regardless of the input context length, with the lowest faithfulness scores mostly observed at the end. Interestingly, we found that the positional discrepancies in generated summaries are not significantly affected by the input context length. Specifically, the decline in faithfulness at the end of summaries remains similar between the 1K and 8K context length settings.
Figure 3: Comparison of faithfulness scores for long summaries generated by different models across increasing output lengths on the Wikipedia dataset, with context lengths ranging from 1K to 8K.
However, the Hallucinate at the Last phenomenon becomes even more severe in long summaries except for Qwen, as shown in Figure 3. For these models, as the length of the generated summary increases, the faithfulness scores toward the end continue to decline, eventually dropping below a faithfulness score of 0.70 when 45 sentences are generated from an 8K context. Qwen, in contrast, often exhibits the largest drop in faithfulness in the middle of the summary, after which the scores recover toward the end.
Tables 3 report the faithfulness and sensitivity of various models. We aggregate the faithfulness scores across context lengths from 1K to 8K and report the overall average. All models except Qwen show markedly low faithfulness scores in the final bin, reflecting high sensitivity. Consistent with the human evaluation in Table 4, Llama demonstrates the highest sensitivity regardless of model size or context length. Qwen, however, is the only model that records sensitivity values below zero. Unlike the consistent decline observed in the other models, Qwen maintains relatively stable faithfulness across bins, resulting in negative sensitivity, and we provide a detailed analysis of these contrasting patterns in Section 4. Supplementary results across context lengths and methods are provided in Appendix C.
The latter part of the summary consistently demonstrates the lowest level of faithfulness across nearly all datasets, with a marked decline once the context length exceeds 7K.
We present the positional faithfulness discrepancy of long summaries generated by the Llama model across different datasets in Figure 4. As the results indicate, the Hallucinate at the Last tendency is consistently observed across multiple domains. Faithfulness scores tend to decrease toward the final segments as the output length increases. In particular, as the length of the generated summary increases, the faithfulness of its later parts continues to decline, with the deterioration becoming especially pronounced once the context length exceeds 7K. Additional sensitivity results across diverse datasets are provided in Appendix D.
@@
Figure 4: Comparison of faithfulness scores in long summaries generated by the Llama3.1-8B-Instruct model across increasing output lengths in multiple domains, with context lengths ranging from 1K to 8K.
@@ 4.1 Is it Intrinsic to Summarization?
We use human-written reference summaries from the CNNDM dataset, selecting examples with long summaries and input contexts of around 2K tokens. We compare these references to summaries generated by the Llama3.1-8B-Instruct model.
@@
Figure 5 compares positional faithfulness between human-written and model-generated CNNDM summaries. We observe that the faithfulness of the generated summaries declines steadily, falling below 0.6 in the final segment, whereas reference summaries dip around the middle but recover. This suggests that Hallucination at the Last is not solely due to summarization’s inherent structure.
@@ 4.2 Is it Intrinsic to Attention?
Unlike previous studies that focus on attention to the first generated token Hsieh et al. (2024), we compute attention weights at the sentence level. Moreover, we segment the full sequence into chunks of 100 tokens. Our analysis specifically focuses on three positions within the generated output: the first sentence, the middle sentence, and the final sentence. Details of the attention weight calculation are provided in Appendix E.
@@
Figure 6 presents a visualization of average attention weights across sentences. As shown in the results, the Llama model, which exhibited the Hallucinate at the Last pattern in Figure 3, assigns nearly three times more attention to the final sentence of the generated summary compared to the first and middle sentences. In contrast, the Qwen model assigns similar levels of attention to all three sentence positions. These findings suggest that increased attention to previously generated text correlates with a higher likelihood of hallucination.
This consistent distribution in Qwen can be attributed to its sliding window attention mechanism Child et al. (2019), which maintains a balanced focus across the context and mitigates the downstream attention concentration that triggers hallucinations. Table 5 compares the Qwen model with and without sliding window attention. In this setting, the sliding window context length is fixed at 16000 tokens. The results clearly show that sliding window attention omission increases sensitivity across all context and summary lengths, up to 9.75. They demonstrate that employing sliding window attention not only enhances decoding efficiency but also reduces attention concentration on later tokens as the generated sequence length increases.
@@
(b) Average attention weights of summaries generated by Qwen.

### v3

@@ Abstract
Large Language Models (LLMs) have significantly advanced text generation capabilities, including tasks like summarization, often producing coherent and fluent outputs. However, faithfulness to the source material remains a significant challenge due to the generation of hallucinations by LLMs. While extensive research focuses on detecting and reducing these inaccuracies, less attention has been paid to the positional distribution of hallucinations within generated text, particularly in long outputs. In this work, we investigate where hallucinations frequently occur in LLM-based long response generation, using long document summarization as a key case study. Focusing on the challenging setting of long context-aware long response generation, we find a consistent and concerning phenomenon: hallucinations tend to concentrate disproportionately in the latter parts of the generated long response. To understand this positional bias, we explore potential contributing factors related to the dynamics of attention and decoding over long sequences. Furthermore, we investigate methods to mitigate this positional hallucination, aiming to improve faithfulness specifically in the concluding segments of long outputs.
@@ 1 Introduction
Figure 1: Comparison of the faithfulness scores of summaries generated around 50, 200, and 800 words using GPT-4o mini.
@@ 2 Generating Summaries
This section describes the experimental setup used to generate the summaries analyzed in our study. We prompt LLMs to generate summaries from original documents. The prompt is provided in Appendix G.
@@ 2.1 Input Context Lengths
We evaluate model performance with input length of over 4K tokens. For the Llama model specifically, we further extend the evaluation to context lengths of 12K and 16K tokens.
@@ 2.2 Output Length Categories
This distinction allows us to compare hallucination patterns in standard summary lengths versus significantly longer generated responses. By including both standard and long output categories, we aim to systematically analyze if and how the positional bias of hallucinations manifests and potentially becomes more severe as the length of the generated summary increases. See Appendix G for more details.
@@ 2.3 Overall Faithfulness
As a preliminary step, we measure the overall faithfulness of the generated long summaries with FineSurE Song et al. (2024) to ensure consistent quality over various context lengths, prior to our main positional analysis. Table 1 shows results for Llama3.1-8B-Instruct Grattafiori et al. (2024) on the Wikipedia dataset, evaluated using FineSurE with GPT-4 Achiam et al. (2023). FineSurE assesses each sentence against the source, identifying error types when inconsistencies occur. The results align with those on CNNDM Hermann et al. (2015), where Llama3-70B-Instruct achieved 85.5% faithfulness. We next conduct a finer-grained positional analysis in Section 3 for various setups.
3 Where Do Hallucinations Most Frequently Occur?
@@
Table 1: Overall faithfulness of long summaries on the Wikipedia dataset, evaluated using FineSurE.
@@ 3.1 Experimental Setup
We evaluate our approach on multiple state-of-the-art LLMs, including Llama3.1-8B-Instruct, Gemma3-12B-Instruct Team et al. (2025), Qwen2.5-7B-Instruct Yang et al. (2024a), and GPT-4o mini, to ensure model-agnostic findings. For Gemma, we set the sliding window attention context length to 1 token (as it cannot be disabled entirely), whereas for Qwen, we modify the default setting of over 131K tokens to 16K tokens. Throughout all experiments, summaries are generated using greedy decoding.
@@
To examine whether the phenomenon is domain-specific, we analyze the positional distribution of faithfulness within generated summaries across various domains. In addition to the Wikipedia dataset used in the main experiment, we include the arXiv Cohan et al. (2018), PubMed Cohan et al. (2018), and GovReport Huang et al. (2021) datasets for further evaluation. For each dataset, we randomly sampled 20 examples for analysis.
@@ 3.2 Evaluation Metric
To enable an in-depth analysis of positional bias within each generated summary, we use the MiniCheck-Flan-T5-Large Tang et al. (2024) model to evaluate faithfulness. MiniCheck is particularly suitable for this fine-grained analysis, as it is an efficient and high-performing fact-checking model that leverages atomic facts Min et al. (2023); Jing et al. (2024); Yu et al. (2024).
Let $s_{i}$ denote the $i$-th sentence in a generated summary $S$. Each sentence is first decomposed into a set of atomic facts $A_{i}$ using an LLM, followed by a filtering step to remove unnecessary atomic facts Yang et al. (2024b), which makes:
@@
We then compare each filtered atomic fact with the source document $D={d_{1},...,d_{M}}$ using the MiniCheck model $\mathcal{M}$ in a pairwise fashion. For each atomic fact, we take the maximum score obtained across all source sentences:
@@
Finally, the sentence-level faithfulness score is computed as the average score of all filtered atomic facts within that sentence:
@@
Each sentence is assigned a faithfulness score in the range $[0,1]$, where 1 indicates that all atomic facts in the sentence are fully supported by the source document, and 0 indicates a complete lack of support.
@@
In addition to the faithfulness score, we propose a simple yet effective metric for analyzing positional discrepancies in generated outputs. Following Singh and Goshtasbpour (2022), each summary is divided into five equal length bins, and the average faithfulness score of the atomic facts within each bin is computed. The bins are defined using the cut function in the pandas package, and the scores of the atomic facts are assigned to bins using open-interval boundaries, meaning that boundary values are not included. Let $f_{k}$ denote the average faithfulness score of the $k$-th bin ($k=1,\dots,5$). We define sensitivity as:
$\text{Sensitivity}=100\times\left(\frac{1}{4}\sum_{k=1}^{4}f_{k}-f_{5}\right).$
A positive value indicates that the faithfulness in the final bin is lower than in the earlier bins (i.e., a tendency to hallucinate at the end), whereas a negative value suggests the opposite trend. Larger sensitivity values correspond to stronger hallucination effects near the end of the generated output.
Additional experimental details and an example of the evaluation procedure are provided in Appendix C. Furthermore, the human evaluation of the quality and completeness of the generated atomic facts and experimental details is presented in Appendix B.
3.3 Hallucinate at the Last
The latter part of the summary exhibits the lowest level of faithfulness across multiple models, consistently observed across summary lengths. The decline becomes most pronounced when the summaries are long.
We present the positional faithfulness discrepancies of both standard and long summaries generated by different models on the Wikipedia dataset in Figure 2. As shown in the figure, both standard and long summaries produced by the three models, excluding Qwen, consistently exhibit a decline in faithfulness toward the later portions of the summary, with the lowest scores typically appearing at the end. This Hallucinate at the Last phenomenon becomes even more pronounced in long summaries: as the summary length increases, the faithfulness near the end continues to decrease, eventually dropping below 0.75 for these models. In contrast, Qwen displays a different pattern. Its faithfulness tends to drop most sharply in the middle of the summary, after which the scores recover toward the end.
Figure 2: Comparison of faithfulness scores for summaries generated by different models on the Wikipedia dataset.
Models Generated Summary Bins Sensitivity
1 2 3 4 5
Llama3.1-8B-Instruct 0.81 0.82 0.79 0.80 0.75 5.50
Llama3.1-70B-Instruct 0.83 0.82 0.83 0.81 0.77 5.25
Gemma3-12B-Instruct 0.85 0.83 0.84 0.83 0.79 4.75
GPT-4o mini 0.83 0.81 0.81 0.78 0.76 4.75
Qwen2.5-7B-Instruct 0.86 0.83 0.83 0.82 0.84 -0.50
Table 2: Faithfulness scores and sensitivity for long summaries generated by various models on the Wikipedia dataset.
@@ 3.3 Hallucinate at the Last
Standard Summary
Llama3.1-8B-Instruct 12K 0.91 0.87 0.85 0.76 0.67 17.65
16K 0.84 0.83 0.76 0.71 0.69 3.00
Long Summary
Llama3.1-8B-Instruct 12K 0.80 0.79 0.66 0.63 0.69 9.50
@@
Table 3: Faithfulness scores and sensitivity for standard and long summaries generated by Llama on the Wikipedia dataset, with 12K and 16K context length.
Model Generated Summary Bins Sensitivity
1 2 3 4 5
Llama3.1-8B-Instruct 0.90 0.87 0.84 0.84 0.70 16.25
@@
Table 4: Human evaluation of faithfulness scores and sensitivity for long summaries generated by Llama on the Wikipedia dataset.
@@
Figure 3: Comparison of faithfulness scores for summaries generated by the Llama3.1-8B-Instruct model in multiple domains.
Table 2 report the faithfulness and sensitivity values for the different models. All models except Qwen show markedly low faithfulness scores in the final bin, indicating high sensitivity. Consistent with the human evaluation in Table 4, Llama demonstrates the highest sensitivity regardless of model size. In contrast, Qwen is the only model that records negative sensitivity values. Unlike the consistent decline observed in the other models, Qwen maintains relatively stable faithfulness across bins, resulting in negative sensitivity. We provide a detailed analysis of these contrasting patterns in Section 4.
Table 3 presents the faithfulness and sensitivity results for both the standard and long summaries generated by the LLaMA model with context lengths of 12K and 16K tokens. Interestingly, we find that positional discrepancies in the generated summaries are not significantly affected by the input context length. As shown in the table, even with longer contexts of approximately 12K and 16K tokens, sensitivity remains high, consistently demonstrating the Hallucinate at the Last phenomenon. Supplementary results across different context lengths and methods are presented in Appendix D.
The latter part of the summary consistently demonstrates the lowest level of faithfulness across all datasets, with a marked decline on the long summaries.
We present the positional faithfulness discrepancy of long summaries generated by the Llama model across different datasets in Figure 3. As the results indicate, the Hallucinate at the Last tendency is consistently observed across multiple domains. Faithfulness scores gradually decrease toward the final segments as the output length increases, and this decline becomes most severe in the concluding portions of longer summaries. Additional sensitivity results across diverse datasets are provided in Appendix E.
The Hallucinate at the Last phenomenon persists across different decoding strategies.
To demonstrate that the Hallucinate at the Last phenomenon is not merely a consequence of error propagation, we first conduct experiments using various decoding strategies. Following  Shi et al. (2024a), we adopt the following decoding strategies to generate summaries.
  • Temperature & Top-p Sampling samples tokens from the next-token probability distribution using the smallest set of tokens whose cumulative probability mass exceeds a given threshold p. We conduct experiments with temperature values $\tau\in\{0.5,0.7\}$ and with top-p thresholds of 0.7 and 0.9.
  • Top-k Sampling samples tokens from the k most probable candidates. We experiment with k values of 20 and 100.
  • $\eta$-Sampling discards tokens whose probabilities fall below an entropy dependent threshold. The hyperparameter $\eta$ is explored with 6e-4 and 4e-3.
We present the positional faithfulness discrepancies in long summaries generated using different decoding strategies in Figure 4. For the Llama and Gemma models, the latter portions of the summaries exhibit the lowest levels of faithfulness across all decoding strategies. This effect is particularly pronounced under $\eta$-sampling for Llama, where the faithfulness score remains above 0.8 in the earlier segments but drops below 0.7 in the final bin. In contrast, the Qwen model shows a different trend. With the exception of two decoding strategies, the faithfulness of the generated summaries does not decline toward the end but instead increases. Notably, the faithfulness scores for Qwen are lowest in the middle portions of the summaries rather than at the end. Motivated by these observations, we further investigate this phenomenon across a broader set of models.
Figure 4: Comparison of faithfulness scores for long summaries generated using different decoding strategies on the Wikipedia dataset.
Human Evaluation
Each atomic fact was labeled as true or false based on whether it was supported by the source document. True labels were scored as 1 and false as 0, and these scores were averaged within the same five positional bins as in the automatic evaluation. The resulting human-judged faithfulness scores are reported in Table 4. Furthermore, to assess the reliability of the human evaluation, we measured inter-annotator agreement among annotators who evaluated the same subset. In total, 543 atomic facts were evaluated, and the annotators agreed on 515 of them, corresponding to a 94.8% raw agreement.
@@ 4.1 Is it Intrinsic to Summarization?
We use human-written reference summaries from the CNNDM dataset, selecting examples with long reference summaries. We compare these references to summaries generated by the Llama3.1-8B-Instruct model.
@@
Figure 5 compares positional faithfulness between human-written and model-generated CNNDM summaries. We observe that the faithfulness of the generated summaries declines steadily, falling below 0.65 in the final segment, whereas reference summaries dip around the middle but recover. This suggests that Hallucination at the Last is not solely due to summarization’s inherent structure.
@@ 4.2 Is it Intrinsic to Attention?
Following Hsieh et al. (2024), we compute attention weights by averaging the decoder self-attention matrices across all layers and heads. Unlike the prior study, however, we compute sentence-level attention by aligning token spans with sentence boundaries and averaging the token-level attention weights within each sentence. We segment the sequence into 100-token chunks and focus our analysis on three positions: the first, middle, and final sentences. Further details are provided in Appendix A.
@@
Figure 6 presents a visualization of average attention weights across sentences. As shown in Figure 6(a), the Llama model, which exhibited the Hallucinate at the Last pattern in Figure 2, assigns nearly three times more attention to the final sentence of the generated summary compared to the first and middle sentences. In contrast, the Qwen model (Figure 6(b)) assigns similar levels of attention to all three sentence positions. These findings suggest that increased attention to previously generated text correlates with a higher likelihood of hallucination.
This consistent distribution in Qwen can be attributed to its sliding window attention mechanism Child et al. (2019), which maintains a balanced focus across the context and mitigates the downstream attention concentration that triggers hallucinations. Table 5 compares Qwen with and without sliding window attention, where the sliding-window context length is fixed at 16,000 tokens. The results clearly indicate that removing the sliding window attention mechanism increases the sensitivity to 3.75 for standard summaries and 3.25 for long summaries. Furthermore, as shown in Figure 6(c), the attention level becomes higher than that observed with sliding window attention. While autoregressive decoders generally allocate more attention to recent tokens, this trend is not uniform across models. Taken together, these results demonstrate that sliding window attention not only enhances decoding efficiency but also reduces attention concentration on later tokens as the generated sequence length increases.
@@
(b) Average attention weights of summaries generated by Qwen with sliding window attention.
(c) Average attention weights of summaries generated by Qwen without sliding window attention.
@@
Models Generated Summary Bins Sensitivity
1 2 3 4 5
Standard Summary
Qwen2.5-7B-Instruct 0.89 0.88 0.85 0.81 0.87 -1.25
w/o sliding window attention 0.90 0.88 0.85 0.83 0.84 2.5(+3.75)
Long Summary
Qwen2.5-7B-Instruct 0.86 0.83 0.83 0.82 0.84 -0.50
w/o sliding window attention 0.86 0.84 0.83 0.82 0.81 2.75(+3.25)